# Volume-Preserving Deformation of Honeycomb Wire Media Enables Broad Plasma Frequency Tunability

### v1

@@ Abstract
Metamaterials are artificially engineered structures designed to possess properties not commonly observed in nature Engheta and Ziolkowski (2006); Kadic et al. (2019); Alù et al. (2024); Sakoda (2019). One notable class of metamaterials, known as wire media, consists of periodically arranged conductive wires embedded in a host medium or in free space Simovski et al. (2012). Wire media are characterized by pronounced spatial dispersion, even at low frequencies Belov et al. (2003); Simovski and Belov (2004), enabling advanced manipulations of electromagnetic fields Simovski et al. (2012); Belov and Silveirinha (2006); Belov et al. (2010); Fernandes, Maslovski, and Silveirinha (2012); Silveirinha and Maslovski (2012).
The interest in wire metamaterials has grown in recent years due to their potential applications in tunable resonators for dark matter searches Kowitt et al. (2023); Lawson et al. (2019); Balafendiev et al. (2022). In particular, the ALPHA consortium employs a wire-medium-filled cavity as a key component of its experimental setup for axion detection – a leading dark matter candidate Millar et al. (2023). A key advantage of wire-medium-filled cavities over conventional designs is their ability to operate at a target resonance frequency largely independent of the cavity volume.
A wire medium Brown (1960); Rotman (1962) refers to a structure composed of parallel metallic wires periodically arranged in a perpendicular plane. This wire metamaterial prevents wave propagation below a characteristic frequency known as the plasma frequency. This frequency depends on the periodicity of the material, i.e. on the geometry of the unit cell Belov, Tretyakov, and Viitanen (2002), on the number of wires and their arrangement within the unit cell Fernandes et al. (2013); Kowitt et al. (2023).
@@
Figure 1: (a) Geometry of a wire metamaterial with a honeycomb lattice of period $a$. A hexagonal unit cell (solid black line) contains parallel wires of radius $r_{0}$. $R$ is the distance between the unit cell center and wires within a unit cell. (b) Breathing deformation of the structure visualization. (c) Numerical calculation of the metamaterial plasma frequency during the breathing deformation. Geometry parameter $R$ changes from $2r_{0}$ to $a/2-r_{0}$ ($a=3$ cm, $r_{0}=0.5$ mm). Insets in the plot show field distributions of the plasma mode for each highlighted $R/a$ value.
Alongside volume preservation, another crucial property for a potential tunable resonator filled with a wire medium is tunability of the working frequency – in this case, the plasma frequency Millar et al. (2023). The tunability in wire media formed by rectangular lattices has been previously investigated in Refs. Belov, Tretyakov, and Viitanen (2002); Kowitt et al. (2023); Sakhno, Balafendiev, and Belov (2024) while other lattices remain unstudied in that context.
To demonstrate a tunability provided by the breathing deformation we performed a numerical simulation in COMSOL COMSOL AB for an infinite wire medium (applying periodic boundary conditions to a hexagonal unit cell) with a period $a=3$ cm composed of infinite and perfectly electric conducting (PEC) wires of radii $r_{0}=0.5$ mm. The plot in Fig. 1(c) shows a dependence of the plasma frequency on a deformation parameter $R$ changing from the minimal possible size equal to $2r_{0}$, where wires touch near the unit cell center, to the maximal $a/2-r_{0}$, where wires from neighboring cells touch near the unit cell boundary.
@@
Table 1: Plasma frequencies obtained from numerical simulations of infinite periodic structure ($a=3$ cm, $r_{0}=0.5$ mm), see Fig. 1 (c).
@@
$\eta_{\text{max}}=\frac{f_{\text{max}}-f_{\text{min}}}{2(f_{\text{max}}+f_{\text{min}})}\times 100\%\approx 78.85\%,$
which significantly surpasses the $\sim 16\%$ Kowitt et al. (2023) and $\sim 26\%$ Balafendiev et al. (2025) tunability reported in previous studies.
@@
The next step of our study was to implement numerical simulations of a finite resonator to assess whether the observed tunability could be realized experimentally. We simulated a hexagonal cavity filled with the metamaterial, as shown in Fig. 2. The hexagonal cavity shape was chosen to achieve optimal termination of the wire medium, enclosing nearly an integer number of hexagonal unit cells. A typical configuration for wire medium cavities involves enclosing the wire medium with metallic walls placed at the edges of the peripheral unit cells. This arrangement supports a fundamental TM mode with a uniform longitudinal electric field and a sinusoidal transverse field profile, and its resonance frequency lies above the plasma frequency Balafendiev et al. (2022). In contrast, an optimized configuration introduces an air gap of quarter-wavelength thickness, $\Delta=\lambda_{p}/4$, between the peripheral unit cells and the cavity walls. This adjustment yields a uniform field distribution for the fundamental TM mode and ensures that its resonance frequency coincides with the plasma frequency Enriquez et al. (2025). In this context, TM modes are defined as those in which the electric field is aligned with the wire axes.
@@
Although the primary goal of this work is to demonstrate a tunable plasma frequency in a wire medium, the wire medium cavities used for experimental validation are also relevant to axion haloscope applications. Simulations indicate that the form factor and quality factor (see Figs. 3(b,c)) of the regular and optimized cavity designs fall within acceptable ranges for such applications Bradley et al. (2003), suggesting that the proposed geometry could serve as a viable platform for future dark matter detection experiments.
@@
To compare with our experimental results, we developed full-wave 3D simulations of the cavities in CST Studio Dassault Systèmes . The simulations included the exact dimensions and material properties of the experimental resonators. Copper was modeled with a conductivity of $\sigma=5.7\times 10^{7}~{}\mathrm{S/m}$, while the FR-4 layers of the PCBs were represented as lossy dielectrics with a relative permittivity of $\varepsilon=4.3$ and a loss tangent of $\tan{\delta}=0.025$.
@@
It is worth mentioning the main sources of discrepancy between our experimental and simulated resonators. On one hand, the excess heat applied during the manual soldering of wires and ports onto the PCBs caused local deformations of the plates, leading to deviations from the ideal behavior assumed in the numerical simulations. In addition, the cavity walls were inserted through the top and bottom plates, requiring a clearance adjustment between the pieces, which introduce variations in the wall spacing $\Delta$. Precise control of this spacing becomes increasingly important at higher frequencies, since the $\lambda_{p}/4$ distance is smaller. This explains why the discrepancy between the simulation and the experiment in Resonator II appears larger than that in Resonator I. The wall spacing directly influences the resonance frequencies of the TM modes in wire-medium-filled resonators, as discussed in Ref. Enriquez et al. (2025).
In conclusion, we numerically and experimentally studied a honeycomb wire metamaterial composed of hexagonal unit cells and demonstrated the significant tunability enabled by a breathing lattice deformation inspired by topological photonics Wu and Hu (2015). To validate the proposed tuning mechanism, we presented numerical results for both infinite wire medium structures and finite wire medium cavities.
@@
While previous methods for tuning the operational frequency of plasma haloscopes achieved tuning ranges below 26% Kowitt et al. (2023); Balafendiev et al. (2025); Millar et al. (2023), the present study numerically and experimentally demonstrates the feasibility of exceeding 64%. Our approach is also competitive with other cavity-based axion search designs: for example, the polygonal coaxial cavity achieves approximately 5% tuning Di Vora et al. (2025), while the thin-shell axion haloscope reports around 12% Dyson et al. (2024). Beyond axion searches, this method constitutes a step forward in structural tunability for metamaterials, where tuning is typically realized by repositioning of the unit cell elements Lapine et al. (2009); Gorkunov and Osipov (2008).
The experimentally observed wide tuning range of the plasma frequency, reaching approximately $64\%$, was achieved without altering the unit-cell volume. This result highlights the potential of honeycomb wire media for designing tunable resonators, particularly since both the form factor and quality factor remain adequate for the intended application. Such resonators are well-suited for future dark matter search experiments Millar et al. (2023), where tuning across a broad frequency range in the GHz regime is essential due to the uncertainty in the axion mass.

### v2

@@ Abstract
Metamaterials are artificially engineered structures designed to possess properties not commonly observed in nature [1, 2, 3, 4, 5]. One notable class of metamaterials, known as wire media, consists of periodically arranged conductive wires embedded in a host medium or in free space [6]. Wire media are characterized by pronounced spatial dispersion, even at low frequencies [7, 8], enabling advanced manipulations of electromagnetic fields [6, 9, 10, 11, 12].
The interest in wire metamaterials has grown in recent years due to their potential applications in tunable resonators for dark matter searches [13, 14, 15]. In particular, the ALPHA consortium employs a wire-medium-filled cavity as a key component of its experimental setup for axion detection – a leading dark matter candidate [16]. A key advantage of wire-medium-filled cavities over conventional designs is their ability to operate at a target resonance frequency largely independent of the cavity volume.
A wire medium [17, 18] refers to a structure composed of parallel metallic wires periodically arranged in a perpendicular plane. This wire metamaterial prevents wave propagation below a characteristic frequency known as the plasma frequency. This frequency depends on the periodicity of the material, i.e. on the geometry of the unit cell [19], on the number of wires and their arrangement within the unit cell [20, 13]. The plasma frequency determines the dispersion of the wire medium [19, 8] and sets the epsilon-near-zero condition [21, 5], which can be tailored to enhance antenna performance [22, 23].
@@
Figure 1: (a) Geometry of a wire metamaterial with a honeycomb lattice of period $a$. A hexagonal unit cell (solid black line) contains parallel wires of radius $r_{0}$. $R$ is the distance between the unit cell center and wires within a unit cell. (b) Breathing deformation of the structure visualization. (c) Numerical calculation of the metamaterial plasma frequency during the breathing deformation. Geometry parameter $R$ changes from $2r_{0}$ to $a/2-r_{0}$ ($a=3$ cm, $r_{0}=0.5$ $m$). Insets in the plot show field distributions of the plasma mode for each highlighted $R/a$ value.
Alongside volume preservation, another crucial property for a potential tunable resonator filled with a wire medium is tunability of the working frequency – in this case, the plasma frequency [16]. The tunability in wire media formed by rectangular lattices has been previously investigated in Refs. [19, 13, 24] while other lattices remain unstudied in that context.
To demonstrate a tunability provided by the breathing deformation we performed a numerical simulation in COMSOL [25] for an infinite wire medium (applying periodic boundary conditions to a hexagonal unit cell) with a period $a=3$ cm composed of infinite and perfectly electric conducting (PEC) wires of radii $r_{0}=0.5$ mm. The plot in Fig. 1(c) shows a dependence of the plasma frequency on a deformation parameter $R$ changing from the minimal possible size equal to $2r_{0}$, where wires touch near the unit cell center, to the maximal $a/2-r_{0}$, where wires from neighboring cells touch near the unit cell boundary.
@@
Table 1: Plasma frequencies obtained from numerical simulations of infinite periodic structure ($a=3$ cm, $r_{0}=0.5$ $m$), see Fig. 1 (c).
@@
$\eta_{\text{max}}=\frac{2(f_{\text{max}}-f_{\text{min}})}{f_{\text{max}}+f_{\text{min}}}\times 100\%\approx 78.85\%,$ (1)
which significantly surpasses the $\sim 16\%$ [13] and $\sim 26\%$ [26] tunability reported in previous studies.
@@
The next step of our study was to implement numerical simulations of a finite resonator to assess whether the observed tunability could be realized experimentally. We simulated a hexagonal cavity filled with the metamaterial, as shown in Fig. 2. The hexagonal cavity shape was chosen to achieve optimal termination of the wire medium, enclosing nearly an integer number of hexagonal unit cells. A typical configuration for wire medium cavities involves enclosing the wire medium with metallic walls placed at the edges of the peripheral unit cells. This arrangement supports a fundamental TM mode with a uniform longitudinal electric field and a sinusoidal transverse field profile, and its resonance frequency lies above the plasma frequency [15]. In contrast, an optimized configuration introduces an air gap of quarter-wavelength thickness, $\Delta=\lambda_{p}/4$, between the peripheral unit cells and the cavity walls. This adjustment yields a uniform field distribution for the fundamental TM mode and ensures that its resonance frequency coincides with the plasma frequency [27]. In this context, TM modes are defined as those in which the electric field is aligned with the wire axes.
@@
Although the primary goal of this work is to demonstrate a tunable plasma frequency in a wire medium, the wire medium cavities used for experimental validation are also relevant to axion haloscope applications. Simulations indicate that the form factor and quality factor (see Figs. 3(b,c)) of the regular and optimized cavity designs fall within acceptable ranges for such applications [28], suggesting that the proposed geometry could serve as a viable platform for future dark matter detection experiments.
@@
To compare with our experimental results, we developed full-wave 3D simulations of the cavities in CST Studio [29]. The simulations included the exact dimensions and material properties of the experimental resonators. Copper was modeled with a conductivity of $\sigma=5.7\times 10^{7}~\mathrm{S/m}$, while the FR-4 layers of the PCBs were represented as lossy dielectrics with a relative permittivity of $\varepsilon=4.3$ and a loss tangent of $\tan{\delta}=0.025$.
@@
It is worth mentioning the main sources of discrepancy between our experimental and simulated resonators. On one hand, the excess heat applied during the manual soldering of wires and ports onto the PCBs caused local deformations of the plates, leading to deviations from the ideal behavior assumed in the numerical simulations. In addition, the cavity walls were inserted through the top and bottom plates, requiring a clearance adjustment between the pieces, which introduce variations in the wall spacing $\Delta$. Precise control of this spacing becomes increasingly important at higher frequencies, since the $\lambda_{p}/4$ distance is smaller. This explains why the discrepancy between the simulation and the experiment in Resonator II appears larger than that in Resonator I. The wall spacing directly influences the resonance frequencies of the TM modes in wire-medium-filled resonators, as discussed in Ref. [27].
In conclusion, we numerically and experimentally studied a honeycomb wire metamaterial composed of hexagonal unit cells and demonstrated the significant tunability enabled by a breathing lattice deformation inspired by topological photonics [30]. To validate the proposed tuning mechanism, we presented numerical results for both infinite wire medium structures and finite wire medium cavities.
@@
While previous methods for tuning the operational frequency of plasma haloscopes achieved tuning ranges below 26% [13, 26, 16], the present study numerically and experimentally demonstrates the feasibility of exceeding 64%. Our approach is also competitive with other cavity-based axion search designs: for example, the polygonal coaxial cavity achieves approximately 5% tuning [31], while the thin-shell axion haloscope reports around 12% [32]. Beyond axion searches, this method constitutes a step forward in structural tunability for metamaterials, where tuning is typically realized by repositioning of the unit cell elements [33, 34].
The experimentally observed wide tuning range of the plasma frequency, reaching approximately $64\%$, was achieved without altering the unit-cell volume. This result highlights the potential of honeycomb wire media for designing tunable resonators, particularly since both the form factor and quality factor remain adequate for the intended application. Such resonators are well-suited for future dark matter search experiments [16], where tuning across a broad frequency range in the GHz regime is essential due to the uncertainty in the axion mass.